# Emergence of Gibbs ensembles as steady states in Lindbladian dynamics

### v2

@@ Abstract
We explicitly construct unique non-equilibrium steady state (NESS) of Lindblad master equation characterized by a Gibbs ensemble $\rho_{\text{NESS}}\propto e^{-\beta\tilde{H}}$, where the effective Hamiltonian $\tilde{H}$ consists only of $U(1)$ conserved charges of the original Hamiltonian. Specifically, when the original Hamiltonian has multiple charges, it is possible to couple them with bathes at different temperature respectively, but still leads to an equilibrium state. Multiple steady states arise if the number of bathes is less than the number of charges. To access the Gibbs NESS, the jump operators need to be properly chosen to fulfill quantum detailed balance condition (qDBC). These jump operators are ladder operators for $\tilde{H}$ and jump process they generate form a vertex-weighted directed acyclic graph (wDAG). By studying the XX model and Fredkin model, we showcase how Gibbs states emerge as steady states.
@@ I Introduction
where $L_{k}$s are jump operators and $H_{\text{eff}}=H-\frac{i}{2}\sum_{k}L_{k}^{\dagger}L_{k}$. The detailed balance condition (DBC) is not required in Lindblad dynamics, making it powerful and capable of characterizing non-equilibrium phenomena [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Recently with development of quantum computation, people regain interests in crafting quantum Gibbs samplers and preparing specific state utilizing Lindblad dynamics [18, 19, 20, 21]. In this work we make direct use of qDBC to derive exact Gibbs state that only depends on $U(1)$ conserved charges of original Hamiltonian.

### v3

@@ Abstract
We explicitly construct unique non-equilibrium steady state (NESS) of Lindblad master equation characterized by a Gibbs ensemble $\rho_{\text{NESS}}\propto e^{-\beta\tilde{H}}$, where the effective hamiltonian $\tilde{H}$ is an element in the center of the commutant algebra $\mathcal{C}$ of the original hamiltonian. Specifically, if $\mathcal{C}$ is Abelian, then $\tilde{H}$ consists only of $U(1)$ conserved charges of the original Hamiltonian. When the original Hamiltonian has multiple charges, it is possible to couple them with bathes at different temperature respectively, but still leads to an equilibrium state. Multiple steady states arise if the number of bathes is less than the number of charges. To access the Gibbs NESS, the jump operators need to be properly chosen to fulfill quantum detailed balance condition (qDBC). These jump operators are ladder operators for $\tilde{H}$ and jump process they generate form a vertex-weighted directed acyclic graph (wDAG). By studying the XX model and Fredkin model, we showcase how Gibbs states emerge as steady states.
@@ I Introduction
where $L_{k}$s are jump operators and $H_{\text{eff}}=H-\frac{i}{2}\sum_{k}L_{k}^{\dagger}L_{k}$. The detailed balance condition (DBC) is not required in Lindblad dynamics, making it powerful and capable of characterizing non-equilibrium phenomena [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Recently with development of quantum computation, people regain interests in crafting quantum Gibbs samplers and preparing specific state utilizing Lindblad dynamics [18, 19, 20, 21]. In this work we make direct use of qDBC to derive exact Gibbs state that only depends on $U(1)$ conserved charges of original Hamiltonian. More generally, the state is in the center of the commutant algebra [22] $\mathcal{C}$ of the original hamiltonian.
@@ II General theory
In summary, we conclude that the steady state is in the center of the commutant algebra $\mathcal{C}$ of the original hamiltonian $H$. The state is by definition in $\mathcal{C}$, and it is also in the bond algebra $\mathcal{A}$ since it is a sum of identities and commutes with every element of $\mathcal{C}$. The double-commutant theorem assures the intersection of $\mathcal{A}$ and $\mathcal{C}$ is the center for both of them [48]. The relevant jump operators lie outside $\mathcal{A}$ and $\mathcal{C}$ for the uniqueness of the steady state, as shown in Fig. 1.
Figure 1: Illustration of the steady state $\rho$, bond algebra $\mathcal{A}$, commutant algebra $\mathcal{C}$ and jump operators $\tilde{L}_{k}$ in the operator space.